OpenStudy (anonymous):

3x + 1 + 5x = 7 +15 + 7x

OpenStudy (anonymous):

8x+1=22+7x x=21

OpenStudy (a1234):

3x + 1 + 5x = 7 +15 + 7x 8x + 1 = 22 + 7x x = 21

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

No problem!